OpenStudy (litzyceron):

Which shows three ordered pairs that are all solutions of the equation y = x + 4? A. (1, 4), (2, 6), and (3, 7) B. (1, 4), (2, 8), and (3, 12) C. (3, 7), (5, 9), and (7, 10) D. (3, 7), (5, 9), and (7, 11)

OpenStudy (litzyceron):

@jtug6

OpenStudy (litzyceron):

@Miguel980

OpenStudy (jtug6):

Answer is D: if y = x+4 try to plug in each X variable to see if its final value is the same as the y variable it's paired with A) (1, 4) x = 1 and 1 + 4 = 5 so if y = 4 in this pair its not equal to 5 so A is a NO B) also has (1, 4) so B is automatically no C) The first two pairs are correct but the last one is (7, 10) and 4+7 = 11 and 11 doesnt equal 10 so C isnt it D) Only last answer but still lets check. Its pair is (7, 11) which matches so it's D

OpenStudy (litzyceron):

thank you again

OpenStudy (litzyceron):

i just have 2 more question more

OpenStudy (jtug6):

Thats fine hopefully you're understanding this. It's nothing more than just plug and chug basically

OpenStudy (litzyceron):

Which three ordered pairs complete the table to give solutions of the equation y = x – 5? A. (–3, –2), (0, –5), and (3, 2) B. (–3, –2), (0, –5), and (3, –2) C. (–3, –8), (0, –5), and (3, 2) D. (–3, –8), (0, –5), and (3, –2)

OpenStudy (jtug6):

Just curious these questions are from algebra 1 I presume?

OpenStudy (litzyceron):

no its from k12 azva

OpenStudy (jtug6):

Oh but I meant is it from what type of math class? Pre-Algebra, Algebra I or Alge II? Or is it different where you're located?

OpenStudy (litzyceron):

ummmm idk ??

OpenStudy (jtug6):

Ok lol. Anyway, it's D

OpenStudy (litzyceron):

ok thank you 1 more

OpenStudy (litzyceron):

Which three ordered pairs complete the table to give solutions of the equation y = –6x? A. (–2, 12), (0,0), and (2, –12) B. (–2, 12), (0,–6), and (2, –12) C. (–2, –12), (0,0), and (2, 12) D. (–2, –12), (0,–6), and (2, 12) https://static.k12.com/calms_media/media/1417500_1418000/1417738/1/d36a266482fe88fd476ffc549481ec30ad544abb/FGA_130919_161121.jpg

OpenStudy (jtug6):

It's A